OpenStudy (anonymous):

∀x,n(x^n∉ℝ) ⇔ x≤0,n∈ℝ,n∉ℚ I asked this question already, but this is, I think, a more correct mathematical statement?

OpenStudy (anonymous):

this is finite math -.-

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

This is innaccurate. I should say ∀x,n(x^n∉ℝ) <- x<0,n∈ℝ,n∉ℚ Finite math?

OpenStudy (anonymous):

I learned this notation in logic theory.

OpenStudy (anonymous):

okies i failed that class

OpenStudy (anonymous):

sorries cant help you there either

OpenStudy (anonymous):

:c it's k

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

\[\forall x,n(x <0,n \in \mathbb{R}, n \notin \mathbb{Q} \rightarrow x^{n} \notin \mathbb{R} )\] is that it? what is it you want to know?

OpenStudy (anonymous):

well, basically x^n, where x<0, and n is a transcendental, x^n MUST be in the complex plane, right? how can we prove this?

OpenStudy (anonymous):

Proof by contrapositive assume \[x \in \mathbb{R} \rightarrow x \ge 0 \cup n \notin \mathbb{R} \cup n \in \mathbb{Q}\]

OpenStudy (anonymous):

Oh ok. Thanks.